OpenStudy (anonymous):

1.80 = x2 / (0.0798-x)

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

is that suppose to be 2x or x^2 ?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

okay multiply both sides by (0.0798-x) to cancel the x in the denominator and get 1.8*(0.0798-x)=x^2 0.14364-1.8x=x^2 move the entire equation to one side 0 = x^2 +1.8x -0.14364 you should recognize what to do at this point if not then let me know and I will help you with the rest

OpenStudy (anonymous):

Ha ok. I forgot to move it all to move side. Been a while since I've had to do something this simple. Thanks